OpenStudy (anonymous):

Solve the equations by substitution. 8x = 2y + 5 3x = y + 7 x = y =

OpenStudy (anonymous):

8x = 2y + 5 3x = y + 7 OR Y=3X-7 8X=2(3X-7)+5 8X=6X-14+5 8X-6X=-9 2X=-9 X=-9/2 X=-4.5 ANS. 8*-4.5=2Y+5

OpenStudy (anonymous):

-36=2Y+5 2Y=-36-5 2Y=-41 Y=-41/2 Y=-20.5 ANS. PROOF: 3*-4.5=-20.5+7 -13.5=-13.5

OpenStudy (anonymous):

The given equations are, 8x = 2y + 5 ...(1) 3x = y + 7 3x-7 = y y = 3x-7 Substituting in (1), we get 3x = 2(3x-7)+5 3x = 6x-14+5 3x-6x = -9 -3x = -9

OpenStudy (anonymous):

So x=-3, y=-9?

OpenStudy (anonymous):

are you sure x=-3?

OpenStudy (anonymous):

-3x=-9 =3 not -3

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

And y=9?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lol..

OpenStudy (anonymous):

Got it, thanks. One problem, I can only give one medal.

OpenStudy (anonymous):

lol well i replied first!

OpenStudy (anonymous):

and i explained it in more detail,

OpenStudy (anonymous):

lool but i am writting whilie giving the anwser i just you are faster than me :P but deserve :) i am new

OpenStudy (anonymous):

lol well i wrote it first, but watever-______-

OpenStudy (anonymous):

lool fine you win :P

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Sorry @nader1 :(

OpenStudy (anonymous):

i just start today to help people :P but anytime for help i am here

OpenStudy (anonymous):

it is kk giver her what she want ^^

OpenStudy (anonymous):

Nice attitude! Btw she can give you a medal.

OpenStudy (anonymous):

if she wantit gona be nice from her if not i respect her :) anyways do you have any question?

OpenStudy (anonymous):

Not right now.